OpenStudy (anonymous):

What is the slope if the line that passes through the pair of points (-7/2, 3) and (-5, 5/2) A. 6/22 B. -6/22 C. 22/6 D. -22/6

OpenStudy (abb0t):

\(\sf slope = \frac{\Delta y}{\Delta x} = \frac{y_2-y_1}{x_2-x_1}=m\)

OpenStudy (abb0t):

All you're doing is taking the difference between each coordinate.